Directions

Precision company wishes to expand but needs a $300,000 loan. The bank requests that Precision prepare a balance sheet and key financial ratios. Precision has kept formal records and is able to provide financial statements as of December 31, 2017. The industry debt ratio averages 45.00%. The industry return on assets is 2.0%. Use the Precision Co. financial statements and Memo template below to complete the requirements.

A. You represent the bank and will present your findings in a Word memo to report the ratios for Precision Co. and identify the conclusions reached from your analysis of the company’s financials. The memo is to be distributed to the VP of the bank, so a well-written and detailed document is crucial, as it will affect the bank leaders’ decision whether or not to lend Precision Co. the $300,000.

B. The body of your memo should be approximately 500 words. Make sure you use complete sentences. Check your work for proper spelling, grammar and punctuation. To receive full credit, you must fully answer each requirement below.

Required:

1a. Compute Precision Co.’s return on total assets for 2017 (Return on total assets is defined in

       Chapter 1).  

1b. Identify and explain its building block category for financial statement analysis.

2a. Compute Precision’s debt ratio for 2017 (Debt ratio is defined in chapter 2).

2b. Identify and explain its building block category for financial statement analysis.

3.    Do you believe the prospects of a $300,000 bank loan are good? Why or why not?

PRECISION COMPANY Comparative Balance Sheets December 31, 2017 and 2016 2017 2016 Assets Current assets 65,000 96,000 ..._._._._...___.... 250,000 265,000 14,000 503,000 Total current assets.... Plant assets 400,000 45,000 350,000 50,000 .. 100,000 100,000 .1.170.000 1.175,000 Liabilities Current liabilities $ 164.000 S 190,000 90,000 12,000 265,000 292,000 Short-term notes payable Taxes payable .. 26,000 Long-term liabilities Notes payable (secured by 665,000 712,000 Stockholders Equity ..544,000 1019,000 Total stockholders equity Total liabilities and equity 966,000 . . ..
Answer #1

1a.> To Calculate Return on Total Assets for 2017 the formula is :

Return on Total Assets = Net Income/Total Assets

ROA = 94000/1684000 = 5.58% (Approx)

1b.> There are 4 Building Block Categories for Financial Statement Analysis.

a) Liquidity and Efficiency

b) Solvency

c) Profitability

d) Market Prospectus

In the given case of Precision Company, the category of Building Block is "Liquidy and Efficiency" as its liqidity ratios i.e. Current ratio and quick ratio is ideal

Current Ratio = Current Assets/Current Liabilities

CA = 514000/265000 = 1.94 : 1 and the ideal Current ratio is 2:1

Further, Quick Ratio = (Current Assets - Inventories)/Current Liabilities

QR = 264000/265000 = 1:1 and the ideal quick ratio is 1:1

Hence, this situation is well suits to Precision Company. Hence, its building block category is "Liquidity and efficiency"   

2.> To Calculate Debt Ratio for 2017 the formula is

Debt Ratio = Total Debt/Total Assets = 501000/1684000 = 0.30

Note : Here totaldDebt equals long term debt and short-term debt. It is not equivalent to total liabilities because it excludes non-debt liabilities such as accounts payable, salaries payable, etc.

3.> We Believe that  prospects of a $300,000 bank loan are good for Precision Company as its the current ratio is near to ideal ratio or we can say that it is ideal ratio . Further, quick ratio of Precision Company is also an ideal ratio. from this precision company will get ability to meet short-term obligations and efficiently generate revenues.